# KP-RED: Exploiting Semantic Keypoints for Joint 3D Shape Retrieval and Deformation

### v2

@@ Abstract
In this paper, we present KP-RED, a unified KeyPoint-driven REtrieval and Deformation framework that takes object scans as input and jointly retrieves and deforms the most geometrically similar CAD models from a pre-processed database to tightly match the target. Unlike existing dense matching based methods that typically struggle with noisy partial scans, we propose to leverage category-consistent sparse keypoints to naturally handle both full and partial object scans. Specifically, we first employ a lightweight retrieval module to establish a keypoint-based embedding space, measuring the similarity among objects by dynamically aggregating deformation-aware local-global features around extracted keypoints. Objects that are close in the embedding space are considered similar in geometry. Then we introduce the neural cage-based deformation module that estimates the influence vector of each keypoint upon cage vertices inside its local support region to control the deformation of the retrieved shape. Extensive experiments on the synthetic dataset PartNet and the real-world dataset Scan2CAD demonstrate that KP-RED surpasses existing state-of-the-art approaches by a large margin. Codes and trained models will be released in https://github.com/lolrudy/KP-RED.
@@ 4.1 Experiments on Full Shapes
To demonstrate the joint R&\&&D ability of KP-RED, we compare our method with the state-of-the-art [41, 21], and present the results in Tab. 1. KP-RED consistently outperforms other competitors in all categories and datasets under the Chamfer Distance metrics, demonstrating our superior ability of R&\&&D under different conditions. In particular, on PartNet dataset, we obtain superior results with a realtive improvement of 85.7%, 86.6% and 84.2% under the average Chamfer Distance, compared with Uy et al. [41] for three categories. When comparing with the second best method ShapeFlow [21], our improvement is still significant with an relative improvement of 73.8%. Moreover, the inference time of Uy et al. [41] reaches 0.7 seconds per instance, while ShapeFlow [21] adopts online optimization for better performance, which leads to non-negligible computational expenses and very long inference time, about 45 seconds per instance. In contrast, KP-RED maintains a real-time inference speed with about 30 ms per instance, whilst surpassing both by a large margin in terms of R&\&&D quality. Fig. 5 demonstrates that our unsupervised keypoint-driven retrieval module is capable of successfully establishing an embedding space for measuring the similarity among objects. As illustrated in Fig. 6, our R&\&&D results are clearly more similar to the target compared to other methods. We attribute this to our well-designed retrieval module, employing local-global feature embedding for effective comparison among objects. Moreover, the cage-based deformation with geometric self-attention preserves structural details and boosts deformation quality. In Supplementary Material, we perform an oracle retrieval experiment to illustrate the effectiveness of our keypoint-guided deformation module.
@@ 4.2 Experiments on Partial Shapes
0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4UCD MetricAverageSF [21]Uy [41]U-RED [11]Ours0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Chair0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Occlussion Ratio [%]UCD MetricTable0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Occlussion Ratio [%]Cabinet

### v3

@@ Abstract
In this paper, we present KP-RED, a unified KeyPoint-driven REtrieval and Deformation framework that takes object scans as input and jointly retrieves and deforms the most geometrically similar CAD models from a pre-processed database to tightly match the target. Unlike existing dense matching based methods that typically struggle with noisy partial scans, we propose to leverage category-consistent sparse keypoints to naturally handle both full and partial object scans. Specifically, we first employ a lightweight retrieval module to establish a keypoint-based embedding space, measuring the similarity among objects by dynamically aggregating deformation-aware local-global features around extracted keypoints. Objects that are close in the embedding space are considered similar in geometry. Then we introduce the neural cage-based deformation module that estimates the influence vector of each keypoint upon cage vertices inside its local support region to control the deformation of the retrieved shape. Extensive experiments on the synthetic dataset PartNet and the real-world dataset Scan2CAD demonstrate that KP-RED surpasses existing state-of-the-art approaches by a large margin. Codes and trained models are released on https://github.com/lolrudy/KP-RED.
@@ 4.1 Experiments on Full Shapes
To demonstrate the joint R&\&&D ability of KP-RED, we compare our method with the state-of-the-art [41, 21], and present the results in Tab. 1. KP-RED consistently outperforms other competitors in all categories and datasets under the Chamfer Distance metrics, demonstrating our superior ability of R&\&&D under different conditions. In particular, on PartNet dataset, we obtain superior results with a relative improvement of 73.8% under the average Chamfer Distance, compared with ShapeFlow [21]. When comparing with the second best method Uy et al. [41], our improvement is still significant with an relative improvement of 37.8%.
Moreover, the inference time of Uy et al. [41] reaches 0.7 seconds per instance, while ShapeFlow [21] adopts online optimization for better performance, which leads to non-negligible computational expenses and very long inference time, about 45 seconds per instance. In contrast, KP-RED maintains a real-time inference speed with about 30 ms per instance, whilst surpassing both by a large margin in terms of R&\&&D quality. Fig. 5 demonstrates that our unsupervised keypoint-driven retrieval module is capable of successfully establishing an embedding space for measuring the similarity among objects. As illustrated in Fig. 6, our R&\&&D results are clearly more similar to the target compared to other methods. We attribute this to our well-designed retrieval module, employing local-global feature embedding for effective comparison among objects. Moreover, the cage-based deformation with geometric self-attention preserves structural details and boosts deformation quality. In Supplementary Material, we perform an oracle retrieval experiment to illustrate the effectiveness of our keypoint-guided deformation module.
@@ 4.2 Experiments on Partial Shapes
0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4AverageSF [21]Uy [41]U-RED [11]Ours0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Chair0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Occlussion Ratio [%]Table0020202020404040406060606080808080000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Occlussion Ratio [%]Cabinet